OpenStudy (anonymous):

A jewelry box has a volume of x3 + 7x2 + 7x – 15 cm3 and the height is x + 3 cm. Find the polynomial that would represent the area of the bottom of the jewelry box? Explain your reasoning.

OpenStudy (anonymous):

V=h*A A=V/h so can you divide x^3 + 7x^2 + 7x – 15 by x+3 ?

OpenStudy (anonymous):

what would the answer be @Algebraic!

OpenStudy (anonymous):

x^3 + 7x^2 + 7x – 15 divided by x+3

OpenStudy (anonymous):

did you try it already?

OpenStudy (anonymous):

yea but i got confused on the sythnic division

OpenStudy (anonymous):